OpenStudy (loser66):

\(\huge Problem1\) We say A is decomposable provided there exist groups B and C, both nonzero, such that A isomorphism to B x C. A nonzero additive abelian group that is not decomposable is said to be indecomposable. We say that A is "uniform" provided every nonzero subgroup of A is indecomposable. Prove that a nonzero additive abelian group is uniform if and only if the intersection of any 2 nonzero subgroups must also be nonzero. \(\huge Problem 2\) 1)Prove that the additive group Z is uniform. 2) Prove that the additive group Q is uniform. 3) Show that the additive group R is not uniform. I do not know how to start. Please help.

OpenStudy (loser66):

My understanding: Let G be an additive abelian group, A1, A2, .....,As are subgroups of G. Since G is abelian, then \(A_i\) is normal. By fundamental theorem of finite abelian group, \(G\simeq A_1\times A_2\times.....\times A_s\) By defining above, each \(A_i\) is indecomposable.

OpenStudy (loser66):

We need show if \(A_i\cap A_j \neq \{0\}\) , then G is uniform. and other direction is G is uniform, then \(A_i\cap A_j \neq \{0\}\) But I don't see how to do. :(

OpenStudy (loser66):

@Alchemista

OpenStudy (loser66):

Give me example, please

OpenStudy (anonymous):

Look at the definition of direct product.

OpenStudy (loser66):

I think of this \(G\simeq Z_2\times Z_3\) and both \(Z_2,Z_3\) are simple, and we know that simple groups are undecomposable. but \(Z_2\cap Z_3 \neq \{0\}\)

OpenStudy (loser66):

I am just down the problem into simple case to understand what is going on and find out how to argue in general.

OpenStudy (loser66):

But Z modulo is so special cases since every element of the set { Z_i} have the same elements. I mean \(Z_2 = \{[0],[1]\}\) , and \(Z_{whatever}=\{ start~from~[0],[1],....~~also\}\) so that intersection of them surely \(\neq \{0\}\)

OpenStudy (anonymous):

Let's try to argue by contradiction.

OpenStudy (anonymous):

Assume that intersection of any two nontrivial subgroups must also be nontrivial. By contradiction, let's suppose that \(A\) is not uniform. In other words there is a nontrivial proper subgroup of \(A\) which is decomposable.

OpenStudy (anonymous):

I think the idea now is to show that if the intersection of every nontrivial subgroup is nontrivial then then the decomposable subgroup cannot exist.

OpenStudy (anonymous):

Look at the section on direct products: http://math.stackexchange.com/questions/106028/semi-direct-v-s-direct-products

OpenStudy (anonymous):

It requires the existence of \(H, K\) where \(H \cap K = \{e\}\) (the intersection is trivial).

OpenStudy (loser66):

We don't have it, right? It asks us to prove that the intersection is not identity.

OpenStudy (anonymous):

Yes I think that's right. I have to check more carefully if it works though.